Gucchi:

Tisha and her academic team are working to go to state finals. They must have a certain number of points, T, to advance. They have had three local matches, a, c and d, and will attend a district match. District match points count for three times the number of points than local matches do. Choose the equation that would help Tisha find how many points they need to earn in the district match, b, to advance. b equals T over 3 minus a minus c minus d b equals T plus a plus c plus d all over 3 b equals T minus a minus c minus d all over 3 b = 3(T − a − b − d)

Gucchi:

Am i right with C?

Gucchi:

@axie

Gucchi:

@carmelle

axie:

what makes you say C?

Gucchi:

I got this equation while reading the question: T = a + c + d + 3b 1. T = a + c + d + 3b First, I subtract a from both sides 2. T - a = c + d + 3b Next, I subtract c from both sides 3. T - a - c = d + 3b Then, I subtract d from both sides 4. T - a - c - d = 3b Finally, I divide both sides by 3 to get C.

Gucchi:

I hope I explained it correctly.

axie:

that math looks exactly like mine, so yes your correct

Gucchi:

thank you